OpenStudy (anonymous):

How many solutions do these equations have? x = y + 1 y = x + 1

OpenStudy (mathteacher1729):

How far have you gotten with this problem so far?

OpenStudy (mathteacher1729):

Hashir -- y = 2x - 3 and y = 4x - 6 are both linear equations. Does that mean they both have only one solution?

OpenStudy (mathteacher1729):

What about y = 2x - 3 and y = 2x - 4. They are both linear equations. Do they also have only one solution? Also, what about y = 2x - 3 and 4y = 8x - 12 ? They both linear equations... do they both only have one solution? :)

OpenStudy (anonymous):

i m talking in general

OpenStudy (anonymous):

and u r posting equation of parallel lines which have no solution

OpenStudy (anonymous):

mind it

OpenStudy (mathteacher1729):

You wrote the number "1" without any explanation beyond "they are both linear equations". This is misleading to a student, because the student could think "Oh, linear equations always have one solution." :) Re-arranging x = y + 1 y = x + 1 the problem is y = x - 1 y = x + 1 Same slope, different intercepts. There are NO solutions to the set of equations given by JP. The answer you gave was incorrect. :(

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

my bad!!!! sorry

OpenStudy (anonymous):

its cool...better than i could have done